# Benchmarking Universal Machine Learning Interatomic Potentials for Elastic Property Prediction

### v1

@@ Abstract
Universal machine learning interatomic potentials have emerged as efficient tool for material simulation fields, yet their reliability for elastic property prediction remains unclear. Here we present a systematic benchmark of four uMLIPs—MatterSim, MACE, SevenNet, and CHGNet—against theoretical data for nearly 11,000 elastically stable materials from the Materials Project database. The results show SevenNet achieves the highest accuracy, MACE and MatterSim balance accuracy with efficiency, while CHGNet performs less effectively overall. This benchmark establishes a framework for guiding model selection and advancing uMLIPs in mechanical property applications.
1School of Intelligence Science and Technology, Nanjing University of Science and Technology, Jiangyin, Jiangsu 214443, China
2School of Physics, Hefei University of Technology, Hefei, Anhui 230009, China
Corresponding author. E-mail: haidi@hfut.edu.cn
1 Introduction
@@
Elastic properties [1], as one of the fundamental properties of materials, play an important role in governing their mechanical behavior across a wide range of applications, from structural engineering to lithium battery systems[2, 3, 4], and other related fields[5]. Accurate prediction of elastic constants and their derived mechanical parameters, such as bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio, is a critical task of computational materials design[6]. Although, the modern density functional theory (DFT) [7, 8, 9, 10]provides reliable and reproducible predictions of elastic properties, they are often associated with heavy computational costs in high-throughput materials screening. Owing to this computational bottleneck, the systematic exploration of large chemical specie spaces is strictly constrained [11], which in turn hinders the efficient evaluation of elastic mechanical properties and delays materials design and discovery.
@@
In recent years, machine learning interatomic potentials (MLIPs)[12, 13, 14, 15, 16] have rapidly emerged as important tools in materials simulation fields, offering an effective balance between the high accuracy of quantum mechanical calculations and the efficiency of classical potentials. Generally, these models are obtained by learning interatomic interactions from large-scale DFT data sets, and enable predictions with near-quantum accuracy while substantially reducing computational cost for crystal structure prediction[17, 18, 2], molecular dynamics simulation [19, 20], and related tasks [21, 22, 23] . Recent advances in graph neural networks, message-passing architectures, and equivariant representations have greatly improved the capabilities of MLIPs. These developments have led to the emergence of universal MLIPs (uMLIPs)[24, 25, 26, 27], which can accurately model a wide range of chemical compositions and crystal structures. However, accurately predicting elastic properties requires a dependable evaluation of the second derivatives of the potential energy surface (PES), which introduces stricter and qualitatively different challenges than those encountered in predicting energies and forces.
@@
So far, many efforts have been paid for developing uMLIPs to improve their accuracy on energies, forces and stress predictions[28]. For instance, the previous research works have shown that uMLIPs on matbench platform[29] perform well in structural optimization, structure prediction, and molecular dynamics simulations tasks. However, their reliability and effectiveness in predicting elastic properties remain unexplored. This because the relationship between energy–force accuracy and second-derivative precision is not straightforward, as elastic constants are highly sensitive to slight variations in the curvature of the PES, which are often difficult to capture with conventional training strategy. Therefore, analyzing the difference between the overall predictive accuracy and property-specific performance of uMLIPs, and evaluating their capability in mechanical property predictions, is of great importance.
@@
In this work, we conduct a systemical evaluation to address the existing gap in crystal mechanical property research.Specifically, we employ four universal machine learning interatomic potentials (uMLIPs) — Crystal Hamiltonian Graph Neural Network (CHGNet)[27], MACE[30], MatterSim[31], and Scalable EquiVariance-Enabled Neural Network (SevenNet)[32] — to calculate the elastic properties of 10,994 crystal structures from the Materials Project database[33, 34, 11, 35, 36, 37, 38], and systematically compare the results with the DFT reference data provided therein. We further quantify model performance differences in key indicators such as shear modulus, bulk modulus, Young’s modulus, Poisson’s ratio, and mechanical stability, as well as computational efficiency. Building on these analyses, we propose evidence-based guidelines for the rational selection of uMLIPs in mechanical property studies.
2 Methodology
2.1 Dataset Construction and Analysis
@@
Figure 1: Crystal structure analysis of the dataset. (a) Periodic table heatmap indicating element occurrence. (b) Sunburst plot illustrating the distribution of crystal systems and space groups, with integers positioned at the margins indicating the corresponding space group numbers. (c) Histogram of number of atoms per unit of cell.
@@
In this work, we collected 10,994 structures with reported elastic properties from the Materials Project database. Among them, 10,871 structures were mechanically stable at the DFT level, and these were used as our benchmark dataset. In the Fig. 1(a), we present the distribution of elements. The nonmetals such as B, C, N, and O, main-group metals like Li and Mg, and transition metals including Ni, Cu, Zn, and Ti appear most often. Heavy and radioactive elements are rare. From crystallographic aspect (see Fig. 1(b)), the dataset covers seven crystal systems. Cubic structures are the most common (23%), followed by tetragonal (20%) and orthorhombic (19%). Trigonal and monoclinic systems make up 16% and 12%, while hexagonal (7%) and triclinic (3%) systems are less frequent. In total, 169 space groups are represented, giving wide crystallographic diversity. Finally, from the number of atoms distribution in Fig. 1(c), we can find that the most structures have fewer than 20 atoms per unit cell, with 5-10 atoms being the most typical, and structures with more than 30 atoms are uncommon.
@@
In the Fig. 2, we present the basic distribution of electronic structure, thermodynamic, and mechanical properties, it can be found that the dataset exhibits a broad and diverse distribution. The statistical analysis shows that 3,248 materials (29.9%) are semiconductors or insulators, with an average band gap of 0.69 eV, while the remaining 7,623 materials (70.1%) are metallic. For the semiconductor subset , as illustrated in 2(a), the majority of structures possess negative formation energies (mean: -0.90 ± 0.98 eV/atom) and energy above hull values (mean: 0.03 ± 0.10 eV/atom ) close to zero, indicating well thermodynamic stability. Regarding mechanical properties 2(b), the dataset shows that the bulk moduli range from 0.33 to 491.33 GPa (mean: 104.41 ± 73.73 GPa), shear moduli from 0.45 to 525.42 GPa (mean: 50.93 ± 44.22 GPa), and Poisson’s ratios from -0.48 to 0.80 (mean: 0.29 ± 0.07). Overall, the dataset demonstrates strong representativeness in electronic, thermodynamic, and mechanical domains, providing a reliable sample for evaluating elastic properties in real materials.
@@
Figure 2: (a) Electronic structure versus thermodynamic stability. Scatter plot of formation energy versus band gap, color-coded by energy above hull. Marginal histograms illustrate the distribution of formation energies and band gaps. (b) Elastic property correlations. Scatter plot of bulk modulus versus shear modulus (log scale), color-coded by Poisson’s ratio. Marginal histograms show the distributions of bulk and shear modulus.
2.2 uMLIP Models Evaluated
@@
In CHGNet[27], the total potential energy is expressed as
@@
MACE[30] advances interatomic potential modeling by combining the systematic completeness of Atomic Cluster Expansion (ACE) with the higher-order equivariant message passing of modern graph neural networks. Unlike conventional message-passing neural networks that primarily encode two-body interactions and rely on deep stacking to capture higher-order correlations, MACE constructs explicit many-body messages within each layer through a hierarchical expansion,
@@
The MatterSim potential[31] is a large-scale, symmetry-preserving machine-learning force field that combines the M3GNet architecture with a periodic-aware Graphormer backbone. Each atomic structure is represented as a graph $G=(Z,V,R,[L,S])$, where atomic nodes $Z$ carry feature vectors $v_{i}$, edges $V$ connect atom pairs $(i,j)$ within a cutoff radius $r_{c}$, $R=\{r_{i}\}$ are atomic coordinates, and $[L,S]$ encodes the global lattice and thermodynamic state. In the M3GNet message-passing block, each edge feature $e_{ij}$ represents the bond between atoms $i$ and $j$, including pairwise information such as chemical type and interatomic distance $r_{ij}=\|r_{i}-r_{j}\|$. To incorporate three-body geometry, $e_{ij}$ is refined through a spherical-Bessel / spherical-harmonic expansion of its neighboring environment:
@@
where $r_{ij}=r_{i}-r_{j}$, $\theta_{jik}$ represents the angle between bonds $e_{ij}$ and $e_{ik}$$j_{\ell}$ and $Y_{\ell}^{0}$ are spherical Bessel functions and spherical harmonics with roots $z_{\ell n}$,
@@
SevenNet[32] (Scalable EquiVariance-Enabled Neural NETwork) follows the atom-decomposed energy formalism widely used in machine-learned interatomic potentials. This locality ensures that the computational cost scales linearly with the number of atoms, $\mathcal{O}(N)$, enabling large-scale molecular dynamics with thousands to millions of atoms. At each message-passing layer $t$, atomic features are updated by
${m}_{v}^{(t+1)}=\sum_{w\in\mathcal{N}(v)}M_{t}\!\big({h}_{v}^{(t)},\,{h}_{w}^{(t)},\,{e}_{vw}^{(t)}\big),\qquad{h}_{v}^{(t+1)}=U_{t}\!\big({h}_{v}^{(t)},\,{m}_{v}^{(t+1)}\big)$ (5)
@@
2.3 Elastic Property Calculations
The second order elastic constant ($C_{ij}$) were calculated using the stress-strain method[39]. According to Hooke’s law, the relationship between stress $\sigma_{ij}$ and strain $\varepsilon_{kl}$ with Voigt notation can be expressed as:
@@ 2.3 Elastic Property Calculations
As for structure optimization and elastic simulation, we use Atomic Simulation Environment (ASE) [40, 41] and Pymatgen [42] softwares. The FIRE algorithm [43] is used for energy minimization and the FretchCellFilter [40] is applied to preserve space group symmetry during relaxation. The Force convergence criteria was set to 0.1 eV/Å for structure relaxation, ensuring mechanical equilibrium before strain application.
Once the elastic tensor is obtained, the bulk modulus, shear modulus, Young’s modulus, Poisson ratio and other derived mechanical properties can be calculated. In this work, all derived mechanical properties are Voigt-Reuss-Hill average values via MechElastic [44] analysis module. The Young’s modulus $E$ and Poisson’s ratio $\nu$ are obtained based on the bulk modulus $K$ and shear modulus $G$ as follows:
@@
3 Results and Analysis
3.1 Model Performance Analysis
@@
Figure 3: Distributions of (a) bulk modulus, (b) shear modulus, (c) Young’s modulus, and (d) Poisson’s ratio, all computed as Voigt–Reuss–Hill (VRH) averages, obtained from DFT and four universal machine-learning interatomic potential models. Each violin shows the full data distribution, where the blue dashed line marks the median, the red dashed line marks the mean, and short lines indicate the minimum and maximum values.
@@
As shown in Figure 3, the distributions of bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio for DFT benchmarks and four uMLIP models are presented, where all values in parentheses in this paragraph denote mean predictions. Overall, all models are able to reproduce the macroscopic trends of DFT, yet systematic deviations remain in the absolute scales.For the bulk modulus, the DFT mean is 104.1 GPa. The mean predictions from all models lie within the range of 100–115 GPa, indicating robust performance in capturing volumetric compressibility. In contrast, the discrepancies are more pronounced for shear and Young’s moduli. The DFT mean values are 47.3 and 122.5 GPa, respectively. CHGNet yields the lowest mean predictions, 28.6 and 77.5 GPa for the shear and Young’s moduli, respectively, systematically underestimating rigidity; MACE (58.2 and 148.1 GPa) and SevenNet (56.3 and 143.9 GPa) both overestimate, reflecting a tendency to over-enhance stiffness; MatterSim (50.8 and 130.5 GPa) gives intermediate results, remaining closest to the DFT benchmarks. For the Poisson’s ratio, the DFT mean is 0.291. CHGNet significantly overestimates (0.371), leading to artificially high ductility predictions; MACE and SevenNet slightly underestimate (0.279 and 0.282), whereas MatterSim (0.294) is nearly identical to DFT. Taken together, these distributional features suggest that while all models capture the overall elastic trends, notable systematic biases remain, particularly in the shear and Young’s moduli as well as in Poisson’s ratio estimation.
@@
Figure 4: (a)-(h) Scatter plot comparison of four uMLIPs against DFT reference values for primary elastic properties: bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio, all in VRH averages. Each subplot shows DFT values on the $x$-axis and ML predictions on the $y$-axis, with the dashed line indicating perfect agreement.
@@
To enable quantitative evaluation, we present point-wise comparisons of the primary elastic properties in Fig. 4. For the bulk modulus, SevenNet and MACE exhibit the highest consistency with DFT, achieving correlation coefficients of approximately $R\approx 0.94$ and mean absolute errors (MAE) around 15 GPa, outperforming both CHGNet ($R=0.909$) and MatterSim ($R=0.924$). For the shear modulus, MACE attains the highest correlation ($R=0.896$), followed by SevenNet ($R=0.895$), while MatterSim yields intermediate accuracy ($R=0.847$) and CHGNet remains significantly weaker ($R=0.546$). Regarding the Young’s modulus, MatterSim yields the mean closest to DFT, but in this correlation-centric assessment MACE attains the higher correlation (R = 0.901) than MatterSim (R = 0.860); SevenNet is lower (R = 0.791), and CHGNet remains the weakest (R = 0.546). CHGNet again shows the weakest performance ($R=0.546$). For the Poisson’s ratio, a different trend emerges: MACE and MatterSim achieve significantly higher correlations ($R\approx 0.65$) than CHGNet ($R=0.301$) and SevenNet ($R=0.374$), indicating their robustness in capturing ratio-type properties. Overall, MACE and SevenNet alternate in leading performance depending on the property considered, while MatterSim also exhibits consistently reliable behavior, achieving mean values closest to DFT and competitive correlations across most properties. The relative superiority of these models remains task-dependent, reflecting the varying accuracy of current universal machine-learning interatomic potentials across different elastic property regimes.
@@
Figure 5: Elastic stability classification analysis comparing DFT and uMLIP models. (a) Stability classification rate showing the percentage of materials predicted as stable by each model. (b) Cross-model stability agreement matrix displaying the percentage of materials with identical stability classifications between each model pair. Stability was determined using either reported elastic stability flags or Born mechanical stability criteria.
@@
Beyond elasticity, stability classification provides another essential benchmark for evaluating model performance. Figure 5 compares the stability predictions of the four uMLIPs against DFT references. SevenNet and MACE achieve the highest performance, with accuracies of 98.3% and 98.1%, respectively, and F1 scores approaching 0.99, reflecting well-balanced capability in identifying both stable and unstable materials. MatterSim ranks closely behind, while CHGNet reaches only 93.4% accuracy, significantly lower than the others and showing a higher rate of missed unstable samples. The confusion matrix analysis further indicates that both MACE and SevenNet exhibit consistently high precision ($\approx 0.997$) and recall (¿0.98), underscoring their robustness and reliability in large-scale stability screening tasks. Cross-model agreement analysis indicates strong overlap, with more than 10,700 materials consistently classified by both MACE and SevenNet in line with DFT. This highlights their superior generalization ability in stability classification across diverse material systems.
@@
3.2 Systematic Error Analysis

### v2

@@ Abstract
Universal machine learning interatomic potentials have emerged as efficient tools for materials simulation, yet their reliability for elastic property prediction remains unclear. Here, we present a systematic benchmark of four uMLIPs—MatterSim, MACE, SevenNet, and CHGNet—against first-principles data for nearly 11 000 elastically stable materials from the Materials Project database. The results show that SevenNet achieves the highest accuracy, MACE and MatterSim balance accuracy with efficiency, while CHGNet performs less effectively overall. To further improve predictive quality, we perform targeted fine-tuning on all four uMLIPs using strained configurations derived from 185 high-error materials. After fine-tuning, CHGNet exhibits the largest overall improvement, with an average mean absolute percentage error reduction of about 23%, followed by MatterSim at around 21% and SevenNet at 18%, whereas MACE shows a performance degradation of roughly 14%. This work provides quantitative guidance for model selection and data refinement, advancing uMLIPs toward reliable applications in mechanical property prediction.
I Introduction
@@
Elastic properties [10.1115/1.3423687], as fundamental properties of materials, play an important role in governing their mechanical behavior across a wide range of applications, from structural engineering to lithium battery systems [crystal3, Kim_2021, XIE2021104602], and related fields [phonons]. Accurate prediction of elastic constants and their derived mechanical parameters, such as bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio, is a critical task of computational materials design [YU2010671]. Although the modern density functional theory (DFT) [hafner2006toward, hohenberg1964inhomogeneous, kohn1965self, curtarolo2005accuracy] provides reliable and reproducible predictions of elastic properties, such calculations are often associated with high computational costs in high-throughput materials screening. Owing to this computational bottleneck, the systematic exploration of large chemical spaces is strictly constrained [de2015charting], which in turn hinders the efficient evaluation of elastic mechanical properties and delays materials design and discovery.
@@
In recent years, machine learning interatomic potentials (MLIPs) [mueller2020machine, friederich2021machine, gokcan2022learning, deringer2019machine, botu2017machine] have rapidly emerged as important tools in materials simulation, offering an effective balance between the high accuracy of quantum mechanical calculations and the efficiency of classical potentials. Generally, these models are obtained by learning interatomic interactions from large-scale DFT data sets, and enable predictions with near-quantum accuracy while substantially reducing computational cost for crystal structure prediction [crystal1, crystal2, crystal3], molecular dynamics simulation [MD1, MD2], and related tasks [related1, related2, cost]. Recent advances in graph neural networks, message-passing architectures, and equivariant representations have greatly improved the capabilities of MLIPs. These developments have led to the emergence of universal MLIPs (uMLIPs) [doi:10.1021/acsami.4c03815, chen2019graph, chen2022universal, deng2023chgnet], which can accurately model a wide range of chemical compositions and crystal structures. However, accurately predicting elastic properties requires a dependable evaluation of the second derivatives of the potential energy surface (PES), which introduces stricter and qualitatively different challenges than those encountered in predicting energies and forces.
@@
So far, many efforts have been devoted to developing uMLIPs that improve the accuracy of energy, force, and stress predictions [deng2025systematic]. For instance, previous studies have shown that uMLIPs on the Matbench platform [dunn2020benchmarking] perform well in structural optimization, structure prediction, and molecular dynamics simulations. However, their reliability and effectiveness in predicting elastic properties remain unexplored. This is because the relationship between energy–force accuracy and second-derivative precision is not straightforward, as elastic constants are highly sensitive to slight variations in the curvature of the PES, which are often difficult to capture with conventional training strategy. Therefore, analyzing the difference between the overall predictive accuracy and property-specific performance of uMLIPs, and evaluating their capability in mechanical property predictions, is of great importance.
@@
In this work, we conduct a systematic evaluation to address the existing gap in crystal mechanical property research. Specifically, we employ four universal machine learning interatomic potentials (uMLIPs) — Crystal Hamiltonian Graph Neural Network (CHGNet) [deng2023chgnet], MACE [batatia2022mace], MatterSim [yang2024mattersim], and Scalable EquiVariance-Enabled Neural Network (SevenNet) [park2024scalable] — to calculate the elastic properties of 10 994 crystal structures from the Materials Project database [horton2025accelerated, jain2013commentary, de2015charting, jain2011formation, wang2021framework, aykol2018thermodynamic, ong2008li], and systematically compare the results with the DFT reference data provided therein. We further quantify model performance differences in key indicators such as shear modulus, bulk modulus, Young’s modulus, Poisson’s ratio, and mechanical stability, as well as computational efficiency. Moreover, we introduce a targeted fine-tuning scheme that augments uMLIP training with strained configurations from high-error materials, enabling a direct evaluation of how incorporating non-equilibrium data affects mechanical predictive accuracy. Building on these analyses, we propose evidence-based guidelines for the suitable selection of uMLIPs and demonstrate that targeted fine-tuning can further enhance their accuracy and reliability in predicting elastic properties.
II Methodology
II.1 Dataset Construction and Analysis
@@
Figure 1: Crystal structure analysis of the dataset. (a) Periodic table heatmap indicating element occurrence. (b) Sunburst plot illustrating the distribution of crystal systems and space groups, with integers placed at the outer margins representing the corresponding space-group numbers. (c) Histogram of atom counts per unit cell.
@@
In this work, we collected 10,994 structures with reported elastic properties from the Materials Project database. Among them, 10,871 structures were mechanically stable at the DFT level, and these were used as our benchmark dataset. In Fig. 1(a), we present the distribution of elements. The nonmetals such as B, C, N, and O, main-group metals like Li and Mg, and transition metals including Ni, Cu, Zn, and Ti appear most often. Heavy and radioactive elements are rare. From a crystallographic perspective (see Fig. 1(b)), the dataset covers seven crystal systems. Cubic structures are the most common (23%), followed by tetragonal (20%) and orthorhombic (19%). Trigonal and monoclinic systems make up 16% and 12%, while hexagonal (7%) and triclinic (3%) systems are less frequent. In total, 169 space groups are represented, giving wide crystallographic diversity. Finally, from the distribution of the number of atoms in Fig. 1(c), we find that most structures have fewer than 20 atoms per unit cell, with 5-10 atoms being the most typical, and structures with more than 30 atoms are uncommon.
@@
In Fig. 2, we present the distribution of the band gap, formation energy, and basic mechanical properties of the dataset, showing that these quantities exhibit a broad and diverse distribution. The statistical analysis shows that 3,248 materials (29.9%) are semiconductors or insulators, with an average band gap of 0.69 eV, while the remaining 7,623 materials (70.1%) are metallic. For the semiconductor subset, as illustrated in Fig. 2(a), the majority of structures possess negative formation energies (mean: -0.90 ± 0.98 eV/atom) and energy above hull values (mean: 0.03 ± 0.10 eV/atom) close to zero, indicating good thermodynamic stability. Regarding mechanical properties in Fig. 2(b), the dataset shows that the bulk moduli range from 0.33 to 491.33 GPa (mean: 104.41 ± 73.73 GPa), shear moduli from 0.45 to 525.42 GPa (mean: 50.93 ± 44.22 GPa), and Poisson’s ratios from -0.48 to 0.80 (mean: 0.29 ± 0.07). Overall, the dataset demonstrates strong representativeness in electronic, thermodynamic, and mechanical domains, providing a reliable sample for evaluating elastic properties in real materials.
@@
Figure 2: Scatter plot of formation energy versus band gap, color-coded by the energy above hull. Marginal histograms illustrate the distribution of formation energies and band gaps. (b) Elastic property correlations. Scatter plot of bulk modulus versus shear modulus (log scale), color-coded by Poisson’s ratio. Marginal histograms show the distributions of bulk and shear moduli.
II.2 Brief Description of Evaluated uMLIPs
@@
In CHGNet [deng2023chgnet], the total potential energy is expressed as
@@
MACE [batatia2022mace] advances interatomic potential modeling by combining the systematic completeness of Atomic Cluster Expansion (ACE) with the higher-order equivariant message passing of modern graph neural networks. Unlike conventional message-passing neural networks that primarily encode two-body interactions and rely on deep stacking to capture higher-order correlations, MACE constructs explicit many-body messages within each layer through a hierarchical expansion,
@@
The MatterSim potential [yang2024mattersim] is a large-scale, symmetry-preserving machine-learning force field that combines the M3GNet architecture with a periodic-aware Graphormer backbone. Each atomic structure is represented as a graph $G=(Z,V,R,[L,S])$, where atomic nodes $Z$ carry feature vectors $v_{i}$, edges $V$ connect atom pairs $(i,j)$ within a cutoff radius $r_{c}$, $R=\{r_{i}\}$ are atomic coordinates, and $[L,S]$ encodes the global lattice and thermodynamic state. In the M3GNet message-passing block, each edge feature $e_{ij}$ represents the bond between atoms $i$ and $j$, including pairwise information such as chemical type and interatomic distance $r_{ij}=\|r_{i}-r_{j}\|$. To incorporate three-body geometry, $e_{ij}$ is refined through a spherical-Bessel / spherical-harmonic expansion of its neighboring environment:
@@
where $r_{ij}=r_{i}-r_{j}$ and $\theta_{jik}$ denotes the angle between bonds $e_{ij}$ and $e_{ik}$; the functions $j_{\ell}$ and $Y_{\ell}^{0}$ correspond to spherical Bessel functions and spherical harmonics with roots $z_{\ell n}$,
@@
SevenNet [park2024scalable] (Scalable EquiVariance-Enabled Neural NETwork) follows the atom-decomposed energy formalism widely used in machine-learned interatomic potentials. This locality ensures that the computational cost scales linearly with the number of atoms, $\mathcal{O}(N)$, enabling large-scale molecular dynamics with thousands to millions of atoms. At each message-passing layer $t$, atomic features are updated by
$m_{v}^{(t+1)}$ $=\sum_{w\in\mathcal{N}(v)}M_{t}\!\big(h_{v}^{(t)},\,h_{w}^{(t)},\,e_{vw}^{(t)}\big),$ (5)
$h_{v}^{(t+1)}$ $=U_{t}\!\big(h_{v}^{(t)},\,m_{v}^{(t+1)}\big).$ (6)
@@
II.3 Elastic Property Calculations
The second-order elastic constants ($C_{ij}$) were calculated using the stress-strain method [TBarron_1965]. According to Hooke’s law, the relationship between stress $\sigma_{ij}$ and strain $\varepsilon_{kl}$ with Voigt notation can be expressed as:
@@ II.3 Elastic Property Calculations
For structure optimization and elastic-property calculations, we employ the Atomic Simulation Environment (ASE) [ase-paper, ISI:000175131400009] and Pymatgen [ong2013python]. The FIRE algorithm [bitzek2006structural] is used for energy minimization and the FretchCellFilter [ase-paper] is applied to preserve space group symmetry during relaxation. The force convergence criterion was set to 0.1 eV/Åfor structure relaxation, ensuring mechanical equilibrium before strain application.
Once the elastic tensor is obtained, the bulk modulus, shear modulus, Young’s modulus, Poisson’s ratio and other derived mechanical properties can be calculated. In this work, all derived mechanical properties are Voigt-Reuss-Hill average values via MechElastic [singh2021mechelastic] analysis module. The Young’s modulus $E$ and Poisson’s ratio $\nu$ are obtained based on the bulk modulus $K$ and shear modulus $G$ as follows:
@@
III Results and Analysis
III.1 Model Performance Analysis
@@
Figure 3: Distributions of (a) bulk modulus, (b) shear modulus, (c) Young’s modulus, and (d) Poisson’s ratio, computed as Voigt–Reuss–Hill (VRH) averages from DFT and the four uMLIP models. Each violin plot shows the overall distribution, with blue and red dashed lines marking the median and mean, and short lines denoting the extrema.
@@
As shown in Fig. 3, we compare the distributions of bulk, shear, and Young’s moduli and Poisson’s ratio obtained from DFT and from the four uMLIP models. Throughout this paragraph, values given in parentheses denote mean model predictions. Overall, all models reproduce the macroscopic DFT trends, but systematic deviations remain in their absolute magnitudes. For the bulk modulus, the DFT mean is 104.1 GPa, while the model means lie in the range 100–115 GPa, indicating robust performance in capturing volumetric compressibility. In contrast, discrepancies are more pronounced for the shear and Young’s moduli: the DFT means are 47.3 and 122.5 GPa, whereas CHGNet yields the lowest values (28.6 and 77.5 GPa), systematically underestimating rigidity; MACE (58.2 and 148.1 GPa) and SevenNet (56.3 and 143.9 GPa) overestimate both moduli, reflecting a tendency to over-enhance stiffness; MatterSim (50.8 and 130.5 GPa) gives intermediate results that remain closest to the DFT benchmarks. For Poisson’s ratio, the DFT mean is 0.291: CHGNet substantially overestimates it (0.371), implying artificially high ductility, MACE and SevenNet slightly underestimate it (0.279 and 0.282), and MatterSim (0.294) is nearly indistinguishable from DFT. Taken together, these distributions indicate that, while all models capture the overall elastic trends, notable systematic biases persist, particularly for the shear and Young’s moduli and for Poisson’s ratio.
@@
Figure 4: (a)–(h) Scatter-plot comparison of four uMLIPs against DFT reference values for primary elastic properties—bulk modulus, shear modulus, Young’s modulus, and Poisson’s ratio, all computed as VRH averages. Each panel shows DFT values on the x-axis and model predictions on the y-axis, with the dashed line representing perfect agreement.
@@
To enable quantitative evaluation, we present point-wise comparisons of the primary elastic properties in Fig. 4. For the bulk modulus, SevenNet and MACE exhibit the highest consistency with DFT, achieving correlation coefficients of approximately $R\approx 0.94$ and mean absolute errors (MAE) around 15 GPa, outperforming both CHGNet ($R=0.909$) and MatterSim ($R=0.924$). For the shear modulus, MACE attains the highest correlation ($R=0.896$), followed by SevenNet ($R=0.895$), while MatterSim yields intermediate accuracy ($R=0.847$) and CHGNet remains significantly weaker ($R=0.546$). Regarding the Young’s modulus, MatterSim yields the mean closest to DFT, but in this correlation-centric assessment MACE attains the higher correlation (R = 0.901) than MatterSim (R = 0.860); SevenNet is lower (R = 0.791), and CHGNet remains the weakest (R = 0.546). For the Poisson’s ratio, a different trend emerges: MACE and MatterSim achieve significantly higher correlations ($R\approx 0.65$) than CHGNet ($R=0.301$) and SevenNet ($R=0.374$), indicating their robustness in capturing ratio-type properties. Overall, MACE and SevenNet alternate in leading performance depending on the property considered, while MatterSim also exhibits consistently reliable behavior, achieving mean values closest to DFT and competitive correlations across most properties. The relative superiority of these models remains task-dependent, reflecting the varying accuracy of current universal machine-learning interatomic potentials across different elastic property regimes.
@@
Figure 5: Elastic stability classification analysis comparing DFT and uMLIP models. (a) Stability classification rate showing the percentage of materials predicted as stable by each model. (b) Pairwise stability-agreement matrix, showing the percentage of materials assigned the same stability outcome across model pairs. Stability was determined using reported elastic stability flags or Born mechanical stability criteria.
@@
Beyond elasticity, stability classification provides another essential benchmark for evaluating model performance. Fig. 5 compares the stability predictions of the four uMLIPs against DFT references. SevenNet and MACE achieve the highest performance, with accuracies of 98.3% and 98.1%, respectively, and F1 scores approaching 0.99, reflecting well-balanced capability in identifying both stable and unstable materials. MatterSim ranks closely behind, while CHGNet reaches only 93.4% accuracy, significantly lower than the others and showing a higher rate of missed unstable samples. The confusion matrix analysis further indicates that both MACE and SevenNet exhibit consistently high precision ($\approx 0.997$) and recall (¿0.98), underscoring their robustness and reliability in large-scale stability screening tasks. Cross-model agreement analysis indicates strong overlap, with more than 10,700 materials consistently classified by both MACE and SevenNet in line with DFT. This highlights their superior generalization ability in stability classification across diverse material systems.
@@
III.2 Systematic Error Analysis